OpenStudy (anonymous):

If f(x) = 3x and g(x) = 3x + 1 + 6. Describe the transformations from f(x) to g(x).

OpenStudy (jdoe0001):

one may note that 1+6 = 7

OpenStudy (jdoe0001):

y = A( Bx + C ) + D D = vertical shift, D > 0, up, D < 0, down